OpenStudy (anonymous):

Solve the system of equations and enter the solution as an ordered pair. 9x - 3y = 63 7x - 3y = 45

OpenStudy (angela210793):

9x - 3y = 63 (- both sides) 7x - 3y = 45 2x=18 x=9 replace x=9 in one of the eq above--->-3y=63-81 y=-6 (18;-6) is the solution